Night Report 2018-11-07

Observers: M. Berton, E. Congiu

Support: S. Benetti, A. Fiore, A. Pastorello

TAT: N. Elias-Rosa, J. Lyman

Standard EFOSC calibrations - biases, dome flats, Rigrz sky flats and arcs (evening and morning)

Weather:

Start - Clear. seeing~1.1

Middle - Clear, seeing~1.2

Issues:

Telescope lost connection for ~1 hr at the very beginning of the night. We started observing at around 1.15 UT

Instrument - EFOSC

UT Target Type Setup airmass seeing notes

---------------------------------------------------------------------------------------

23.50 Feige110 STD all 1.13 1.1 Telescope connection went down. We'll take the standard in the morning.

01.00 Image Analysis

01.18 SN2018coq follow gr13 1.3 0.8

02.25 SN2009ip follow BVRi 1.2 0.7 telescope focus at the beginning

04.01 SN2018gft follow gr11 1.3 0.8

04.53 SN2018gft follow gr16 1.5 1.2

05.53 SN2018hti follow gr13 1.4 1.2

07.00 AT2018hwp class gr13 1.4 1.0 galaxy

07.20 SN2018hjw follow gr13 1.5 1.1

07.50 AT2018hwm class gr13 1.4 0.8 Type II?

08.08 AT2018hyw class gr13 1.8 1.0 Type IIP

08.25 LTT3864 STD all 1.4 0.9